# Leveraging Synthetic Data for Question Answering with Multilingual LLMs in the Agricultural Domain

### v1

@@ Abstract
Enabling farmers to access accurate agriculture-related information in their native languages in a timely manner is crucial for the success of the agriculture field. Although large language models (LLMs) can be used to implement Question Answering (QA) systems, simply using publicly available general-purpose LLMs in agriculture typically offer generic advisories, lacking precision in local and multilingual contexts due to insufficient domain-specific training and scarcity of high-quality, region-specific datasets. Our study addresses these limitations by generating multilingual synthetic agricultural datasets (English, Hindi, Punjabi) from agriculture-specific documents and fine-tuning language-specific LLMs. Our evaluation on curated multilingual datasets demonstrates significant improvements in factual accuracy, relevance, and agricultural consensus for the fine-tuned models compared to their baseline counterparts. These results highlight the efficacy of synthetic data-driven, language-specific fine-tuning as an effective strategy to improve the performance of LLMs in agriculture, especially in multilingual and low-resource settings. By enabling more accurate and localized agricultural advisory services, this study provides a meaningful step toward bridging the knowledge gap in AI-driven agricultural solutions for diverse linguistic communities.
@@
Rishemjit Kaur1,2,*, Arshdeep Singh Bhankhar1,*, Surangika Ranathunga3,*, Jashanpreet Singh Salh1, Sudhir Rajput1, Vidhi1, Kashish Mahendra1, Bhavika Berwal1, Ritesh Kumar 1,2, 1CSIR-Central Scientific Instruments Organisation, India 2Academy of Scientific and Innovative Research (AcSIR), Ghaziabad, U.P., India 3School of Mathematical and Computational Sciences, Massey University, New Zealand * These authors contributed equally Correspondence: rishemjit.kaur@csio.res.in, s.ranathunga@massey.ac.nz
@@ 1 Introduction
Agriculture forms the backbone of many economies, specially for countries such as India, with agriculture being the primary occupation for a major part of the population (Ministry of Agriculture & Farmers Welfare, Government of India, 2024). Despite substantial advancements of Artificial Intelligence (AI) across healthcare, finance, and education (Li et al., 2024; Zhu et al., 2024), the agriculture sector has seen relatively limited integration of cutting-edge technologies such as Large Language Models (LLMs). Experiments with general-purpose LLMs in agricultural settings suggest that these models can provide useful assistance but often default to generic advice, lacking in local specificity Ibrahim et al. (2023); Calone et al. (2024). Agriculture-specific chat-bot systems such as Farmer.Chat Singh et al. (2024) that have multilingual support are promising, however there has been no quantitative evaluation of such systems. On the other hand, the handful of Agriculture-specific Question Answer (QA) datasets are only for high resource languages such as English, German and Chinese Dofitas et al. (2025); Jiang et al. (2025); Kästing and Hänig (2024); Kpodo et al. (2024).
In this paper, we present a novel synthetic QA dataset in three languages - English, Hindi and Punjabi, for the agriculture domain. In addition, we provide human-curated evaluation datasets for all three languages. We carry out extensive evaluations to determine the best LLM, as well as the prompt. In contrast to previous research that used either metrics such as BLUE and ROGUE Johnny and Jaya Nirmala (2022); Jiang et al. (2025) or LLM-as-judge Yang et al. (2024); Gauba et al. (2025) approaches for evaluation, we employ human evaluation at each step. Through this, we highlight the inadequacy of automated evaluation schemes for domain, as well as country-specific QA evaluation tasks. With the synthetic training data, we carry out two different experiments (i) train and test LLMs with language-specific data (ii) train with English data and use translate-test approachAhuja et al. (2023), where the questions in Hindi or Punjabi were first translated into English, fed to English trained model, and the generated responses were then translated back into Hindi or Punjabi, respectively. Our results demonstrate notable improvements from language-specific fine-tuning and further gains using the translate-test approach via an instruction-tuned English model. In Punjabi, factuality improved by 7.4% with fine-tuning and further increased by 8.7% using the translate-test approach. In Hindi, factuality rose by 10.8% through fine-tuning and gained an additional 10.1% with the translate-test method. The English model also demonstrated substantial improvements, with factuality increasing by 11.5%, and relevancy rising by 25.2% after fine-tuning.
These consistent improvements across all metrics underscore two key insights: (1) domain-specific fine-tuning significantly boosts performance, and (2) instruction-tuned English models, when used with translate test approach, outperform native language fine-tuned models
@@ 2.1.1 Data Source
The agricultural data used to generate QA pairs was collected through web scraping from Vikaspedia [^1] https://en.vikaspedia.in/, an online encyclopedia created by the Government of India. A detailed summary of data source files and token statistics across these domains is presented in Table 1222Token counts computed using the LLaMA 3 8B Instruct tokenizer..
@@
The raw text in each file was split into manageable chunks of 2000 characters each, with an overlap of 200 characters between consecutive chunks. These chunk and overlap sizes were chosen based on experimentation to ensure that important contextual information was retained across chunks.
Instruction:
Suppose you are a knowledgeable expert in agriculture.
No additional context is required.
Context:
{context}
Response:
Question: [Question here]
Answer: [Answer here]
Examples:
{examples}
Figure 1: Prompt used for synthetic data generation. When prompting the model, the context tag will be replaced with a fixed input chunk from the agricultural dataset.
2.1.2 QA Generation Using LLMs
Paula et al. (2024) demonstrated that QA generation by prompting an LLM in a few-shot setting is more effective than approaches such as end-to-end generation (the sequence of questions and answers is generated given the input context) and pipeline methods (the answers are extracted first and questions are then generated in an answer-aware manner) and was therefore used for QA generation in our work.
A total of 60,140 question-answer pairs were created from the agricultural dataset in English. This data set was further translated into Hindi and Punjabi, using IndicTrans2 Gala et al. (2023).
Model Question and Answer
LLaMA Q: How can I manage thrips in groundnut?
A: Apply Fipronil 2 ml/litre.
Table 2: Comparison of a single QA pair generated by LLaMA and DeepSeek models.
2.2 Evaluation Dataset
We manually created an English test dataset by collecting QA pairs from multiple sources such as Tamil Nadu Agricultural University [^3] http://www.agritech.tnau.ac.in/expert_system/, Union Bank of India-FAQ-Agri-Products [^4] https://www.unionbankofindia.co.in/pdf/faq-agri-products.pdf, ICAR-NRRI, Cuttack [^5] https://icar-nrri.in/faq/, and ICAR-NRCP, Solapur [^6] https://nrcpomegranate.icar.gov.in/FAQ. From these sources, we manually generated 1,699 QA pairs, ensuring diversity in topics, relevance to agricultural advisory, and high-quality reference answers.
Additionally, a Hindi test dataset was developed by transcribing videos from the DD Kisan youTube channel [^7] https://www.youtube.com/watch?v=MYkBswpxRcA. 50 videos were selected at random to cover diverse agricultural topics. GenAI 2.0 Flash  Google DeepMind (2025) was employed to generate QA pairs from 500 token segments of the transcript, resulting in 1,200 Hindi QA pairs. Subsequently, these were translated into Punjabi using IndicTrans2 Gala et al. (2023). The Hindi and Punjabi data sets were manually cleaned and verified to ensure an accurate and high quality test data set.
3 Implementation Details
3.1 Experiments
Selection of LLMs: The LLMs selected for initial zero-shot experiments are shown in Table 3.
Punjabi Hindi English
Llama 3.1 Llama 3.1 Airavata
Llama 3 Llama 3 Llama 3
Navarasa Navarasa Deepseek
Mistral Airavata
Gemma
Table 3: Models selected for each language. The best performing models for each language were fine-tuned and are highlighted in bold.
Selection of Prompt: We started with the prompt provided by Silva et al. (2023) which was originally designed for agriculture examination contexts. However, our long-term goal is to support farmers with advisories. Hence, we adapted the prompt and generated six different prompt variations (see Appendix A) . The intuition behind each prompt is given in the Appendix B. Zero-shot experiments were carried out to select the best prompt which was used for further fine-tuning of models in all languages.
LLM Fine-tuning: The best performing model under the zero-shot setting was fine-tuned with the synthetic dataset we created. Further implementation details are provided in Appendix F, and the hyperparameters are listed in Appendix E. Fine-tuned models for each language were evaluated using their respective language-specific evaluation datasets. Further, we also conducted an experiment to evaluate the efficacy of the translate-test approach for Hindi and Punjabi languages.
3.2 Evaluation Metrics
The responses were evaluated by human evaluators using three main metrics: Agricultural Consensus, Relevancy, and Factuality. These metrics are inspired by similar evaluation methods used in prior research Singhal et al. (2023); Balaguer et al. (2024). More details about these metrics are provided in Table 4. Human evaluators followed evaluation guidelines as outlined in Appendix  C. Evaluators are native speakers of Hindi or Punjabi and are also fluent in English.
Multiple raters/measurements Koo et al. (2016) was used to calculate inter-annotator agreement scores for relevancy and factuality. Fleiss Kappa  Elangovan et al. (2025) was used for reliability. Justification for the selection of these measures and how to interpret the results are discussed in Appendix H.
No. Parameter Evaluation Question Scale
Table 4: Evaluation criteria for human assessment
4 Results
Due to the scarcity of human evaluators, it was not possible to evaluate all the LLMs across all the data samples in the test set. Therefore, a subset of QA pairs was selected. From each QA pair, the question was used in the prompt, and the corresponding answer generated by the LLM was evaluated against the ground-truth answer by a human. Table 5 shows the amount of human evaluations carried out for the purpose of LLM selection.
4.1 Model and Prompt Selection
Setting Models/ Temp Unique Total
Prompts QA pairs Evals
Multilingual Model Comparison
English 4 LLMs 3 20 720*
Hindi 5 LLMs 1 100 1500*
Punjabi 3 LLMs 1 100 900*
Prompt Selection (1 LLM)
English 6 Prompts 1 25 150
Base vs. Finetuned Evaluation
English 2 (Base+FT-EN) 1 300 600
Hindi 2 (Base+FT-HI) 1 240 480
Punjabi 2 (Base+FT-PA) 1 240 480
Translate Test Approach (via FT-EN model)
Hindi 1 (FT-EN) 1 240 240
Punjabi 1 (FT-EN) 1 240 240
Total Evaluations 5,310
Table 5: Overview of all experiments. Three temperature values (0.2, 0.5, 0.7) were tested only for English model selection. The best temperature identified in that step was used for other experiments.
*Each QA pair in the multilingual model comparison experiment was independently evaluated by 3 annotators. All other evaluations were conducted by a single annotator.
77footnotetext: *Each QA pair in the multilingual model comparison experiment was independently evaluated by 3 annotators.
The results in Table 6 show that LLaMA-3-8B-Instruct, LLaMA-3.1-8B-Instruct and LLaMA-3.1-8B-Instruct emerged as the best performing models for English, Hindi and Punjabi, respectively, at temperature T = 0.2. The experiments were also conducted at T =[0.5, 0.7] and results are in Appendix I. Prompt 1 (see Appendix A)was used for these experiments. Inter-rater agreement was found to be $>$ 0.62 for all cases showing substantial agreement among annotators (see Table 7) .
Language Model Name Consensus Relevancy Factuality
English Airavata-7B-Instruct 100% 58.3% 30.0%
DeepSeek-8B-Instruct 100% 33.3% 28.3%
LLaMA-3-8B-Instruct 100% 98.3% 66.6%
Mistral-7B-Instruct 98.3% 93.3% 30.0%
Hindi Gemma-7B-Instruct 90% 26.3% 14.6%
LLaMA-3-8B-Instruct 84% 25.3% 13.3%
LLaMA-3.1-8B-Instruct 96% 74.6% 35.6%
Mistral-7B-Instruct 100.0% 37.3% 5.6%
Navarasa-7B-Instruct 91.0% 58.6% 21%
Punjabi LLaMA-3-8B-Instruct 87% 17.6% 12.3%
LLaMA-3.1-8B-Instruct 100% 43.3% 24.3%
Navarasa-7B-Instruct 95.6% 29.3% 15.6%
Among the six tested prompts, Prompt 1 achieved the highest scores across all evaluation metrics and was therefore selected as the final instruction format. The evaluation results are shown in Table 8.
Metric Calculation English Hindi Punjabi
Consensus Fleiss’ $\kappa$ 1.0 1.0 0.95
Relevancy ICC 0.79 0.62 0.77
Factuality ICC 0.82 0.93 0.75
Table 7: Inter-rater agreement for three languages across evaluation criteria.
4.2 Base vs. Fine-Tuned Model Results
For each language, the model that provided the best result in the zero-shot setup was used for fine-tuning experiments. Results are reported in Table 9. A notable improvement was observed for relevancy and factuality across all three languages after fine-tuning, despite a drop in performance for consensus for Hindi and Punjabi. The results in Table 9 indicated that the English fine-tuned model in translate-test approach outperformed the Hindi and Punjabi fine-tuned models.
Prompt Number Consensus Relevancy Factuality
Prompt 1 89.3% 90.6% 70.6%
Prompt 2 89.3% 80.0% 65.3%
Prompt 3 89.3% 77.3% 66.6%
Prompt 4 89.3% 54.6% 58.6%
Prompt 5 89.3% 74.6% 60.0%
Prompt 6 89.3% 80.0% 58.6%
Table 8: Human evaluation results for the selected English prompts.
Lang Model Consensus Relevancy Factuality
BASE FT TT BASE FT TT BASE FT TT
En LLaMA-3 96.0% 97.7% 59.0% 84.7% 46.6% 58.1%
Hi LLaMA-3.1 98.5% 96.7% 97.92 64.7% 67.4% 71.53 28.1% 38.9% 49.03
Pa LLaMA-3.1 97.4% 94.5% 98.75 41.0% 61.3% 73.47 21.4% 28.8% 37.50
Table 9: Comparison of Base, Fine-Tuned (FT) and Translate-test with En-FT model (TT) Performance Across Languages
5 LLM-as-a-Judge Evaluation
In the above experiments, 5,310 human evaluations were done. Although human evaluations are the gold standard for assessing model performance, they can be time consuming, expensive, and prone to differences in judgment between reviewers  Tan et al. (2025). To find more scalable alternatives, we employed the "LLM-as-a-Judge" approach for evaluation. We used three instruction-tuned models, namely Llama-4-Scout-17B-16E-Instruct, Llama-4-Maverick-17B-128E-Instruct, and DeepSeek-R1-Distill-Llama-70B [^8] We excluded larger models like Llama 4 Behemoth (288B) and DeepSeek-R1 (685B) due to high computational costs and resource constraints.. We generated prompts, each corresponding to the evaluation metrics ie. consensus, factuality and relevancy. See Appendix  A for prompts), which were adapted from G-eval  (Liu et al., 2023). The same QA pairs used for human evaluation were also evaluated using LLM as the judge, across all languages.
For the relevance and factuality metrics, the Spearman correlation  Spearman et al. (1904) was used to assess the agreement between the LLM’s judgments and those of human evaluators. For the agriculture consensus metric, which involves binary classification, refer to Appendix J. The mean rating from multiple evaluators was computed for each QA pair, and this average score was then correlated with the LLM’s corresponding rating. Llama-4-Scout-17B-16E-Instruct was identified as the best-performing model based on correlation scores.
Lang. Model Rel. (Spearman $\rho$) Fact. (Spearman $\rho$)
En DeepSeek 0.25 0.10
L4-Scout 0.27 0.32
L4-Maverick 0.21 0.18
Hi DeepSeek 0.54 0.47
L4-Scout 0.50 0.53
L4-Maverick 0.45 0.46
Pa DeepSeek 0.33 0.16
L4-Scout 0.39 0.34
L4-Maverick 0.28 0.38
Table 10: Correlation between LLM judges and human evaluators across languages and metrics. Relevancy and Factuality use Spearman’s $\rho$.
6 Conclusion
In this work, we introduced a multilingual fine-tuning dataset in English, Hindi, and Punjabi for Indian agricultural context, addressing the need for regionally and linguistically relevant datasets in this domain. We also proposed an evaluation dataset that can serve as a benchmark for assessing the performance of LLMs on agricultural queries. Our experiments demonstrate that fine-tuned models significantly outperform their base counterparts. Furthermore, for Hindi and Punjabi queries, a translate-test approach where the queries are translated into English and processed using the English fine-tuned model, yields better performance than directly fine-tuning models in those languages. These findings underscore the potential of cross-lingual transfer and resource-efficient strategies for building multilingual agricultural advisory systems. As the next step, we plan to validate our models and system through field testing with farmers.
7 Limitations
There are several limitations to our work. First, the data set created and the methodology used in this study include only textual data. However, in real-world scenarios, farmers may have image-based queries such as pictures of crops or pests/diseases. It is not addressed by our current approach.Our future work will focus on building a multimodal and multilingual dataset that can handle both textual and visual input. It will also incorporate a wider range of Indian languages.Second, human evaluation was done on a subset of the dataset, due to the resource-intensive nature of the task. In future, a more comprehensive human evaluation across the entire dataset is necessary to ensure model’s efficacy and deployment. Another limitation is that is that our experiments were conducted using relatively smaller models (7B to 8B parameter range) due to limited computational resources. In future, the effect of using larger models on performance can be explored. Our current work does not account for dynamic contextual factors such as weather conditions or soil health, which are often important in generating accurate and dynamic advisories. Integrating such contextual data through various APIs in an agentic framework can enable adaptive and real-time responses and it is an important direction for our future work. The human annotators volunteered their time to contribute to the project, and no financial compensation was provided.
Appendix A Tested Prompts
This section provides the six tested prompts used for evaluating Mistral-7B in the Prompt Selection experiment. Each prompt was designed to optimize the quality of the response for agricultural question-answer generation.

### v2

@@ Abstract
Enabling farmers to access accurate agriculture-related information in their native languages in a timely manner is crucial for the success of the agriculture field. Publicly available general-purpose Large Language Models (LLMs) typically offer generic agriculture advisories, lacking precision in local and multilingual contexts. Our study addresses this limitation by generating multilingual (English, Hindi, Punjabi) synthetic datasets from agriculture-specific documents from India and fine-tuning LLMs for the task of question answering (QA). Evaluation on human-created datasets demonstrates significant improvements in factuality, relevance, and agricultural consensus for the fine-tuned LLMs compared to the baseline counterparts.
@@
Rishemjit Kaur1,2,*, Arshdeep Singh Bhankhar1,*, Jashanpreet Singh Salh1, Sudhir Rajput1, Vidhi1, Kashish Mahendra1, Bhavika Berwal1, Ritesh Kumar 1,2, Surangika Ranathunga3,*, 1CSIR-Central Scientific Instruments Organisation, India 2Academy of Scientific and Innovative Research (AcSIR), Ghaziabad, U.P, India 3School of Mathematical and Computational Sciences, Massey University, New Zealand * These authors contributed equally Correspondence: rishemjit.kaur@csio.res.in, s.ranathunga@massey.ac.nz
@@ 1 Introduction
Agriculture forms the backbone of many economies, specially for countries such as India. Despite substantial advancements of Artificial Intelligence across areas such as healthcare and education (Zhu et al., 2024), agriculture sector has seen relatively limited integration of cutting-edge technologies such as Large Language Models (LLMs). General-purpose LLMs have been shown to provide useful assistance in agricultural settings, but their advice is often generic, lacking in local specificity Ibrahim et al. (2023); Calone et al. (2024). Agriculture-specific chat-bot systems such as Farmer.Chat Singh et al. (2024) that have multilingual support are promising, however there has been no quantitative evaluation of such systems. On the other hand, the handful of agriculture-specific Question Answer (QA) datasets are only for high resource languages such as English, German and Chinese Dofitas et al. (2025); Jiang et al. (2025); Kästing and Hänig (2024).
In this paper, we present a novel synthetic QA dataset in three languages - English, Hindi and Punjabi, for the agriculture domain. In addition, we provide human-curated evaluation datasets for all three languages. We carry out extensive human evaluations to determine the best LLM, as well as the prompt. Contrary to prior research that used either metrics such as ROGUE Johnny and Jaya Nirmala (2022); Jiang et al. (2025) or LLM-as-judge Yang et al. (2024); Gauba et al. (2025) for agriculture QA evaluation, we employ human evaluation at each step. Through this, we highlight the inadequacy of automated evaluation schemes for domain- and country-specific QA evaluation tasks.
With the synthetic training data, we carry out two different fine-tuning experiments (i) train and test LLMs with language-specific data (ii) train with English data and use translate-test approach Ahuja et al. (2023), where the questions in Hindi/Punjabi are first translated to English, fed to the LLM trained with English data, and translating the generated responses back to Hindi/Punjabi.
Although the Llama-3x instruct models showed over 95% for consensus, relevancy and factuality results were rather low, with the average across languages being 44.3% and 29%, respectively. However, after fine-tuning with synthetic datasets, these averages increased by 22.1% and 14% (respectively). The translate-test approach further increased results for Hindi and Punjabi. These improvements underscore two key insights: (1) domain-specific fine-tuning significantly boosts performance, and (2) instruction-tuned English LLMs, when used with translate test approach, outperform native language fine-tuned LLMs. Our data (under CC license) and models will be made public .
@@ 2.1 Synthetic Training Data Generation
The agricultural data used to generate QA pairs was collected through web scraping from Vikaspedia [^1] https://en.vikaspedia.in/, an online encyclopedia created by the Government of India. It contains data across diverse topics such as fisheries and women in agriculture. agriculture are also documented. A summary of data source files and token statistics across these topics is presented in Table 10 in Appendix A. The raw text was split into chunks of 2000 characters each, with an overlap of 200 characters between consecutive chunks. These chunk and overlap sizes were chosen upon experimentation.
Instruction:
Suppose you are a knowledgeable expert in agriculture.
No additional context is required.
Context:
{context}
Response:
Question: [Question here]
Answer: [Answer here]
Examples:
{examples}
Figure 1: Prompt used for synthetic data generation. During prompting, the context tag is replaced with a fixed input chunk from the agricultural dataset.
Paula et al. (2024) showed that QA generation by prompting an LLM in a few-shot setting is more effective than approaches such as end-to-end generation (sequence of questions and answers is generated given the input context) and pipeline methods (answers are extracted first and questions are then generated in an answer-aware manner) and was therefore used for QA generation in our work.
A total of 60,130 question-answer pairs were created from the agricultural dataset in English. This dataset was further translated into Hindi and Punjabi, using IndicTrans2 Gala et al. (2023).
Model Question and Answer
Llama Q: How can I manage thrips in groundnut?
A: Apply Fipronil 2 ml/litre.
Table 1: Comparison of a single QA pair generated by Llama 3.1 and DeepSeek models.
2.2 Evaluation Dataset
We manually created an English test dataset by collecting 1674 QA pairs from multiple sources such as Tamil Nadu Agricultural University, Union Bank of India-FAQ-Agri-Products, ICAR-NRRI, Cuttack, and ICAR-NRCP, Solapur. The links to sources are provided in Appendix B.
Additionally, a Hindi test dataset was developed by transcribing 50 videos from the DD Kisan youTube channel [^2] https://www.youtube.com/watch?v=MYkBswpxRcA. GenAI 2.0 Flash  Google DeepMind (2025) was employed to generate QA pairs from 500 token segments of the transcript, resulting in 1,219 Hindi QA pairs. Subsequently, these were translated into Punjabi using IndicTrans2 Gala et al. (2023). The Hindi and Punjabi datasets were manually cleaned and verified to ensure an accurate and high quality test dataset.
3 Implementation Details
3.1 Experiments
Selection of LLMs: The LLMs selected for initial zero-shot experiments are shown in Table 2. (Further details of LLMs are in Appendix C).
Punjabi Hindi English
Llama 3.1 8B Llama 3.1 8B Airavata 7B
Llama 3 8B Llama 3 8B Llama 3 8B
Navarasa 7B Navarasa 7B Deepseek 8B
Mistral 7B Mistral 7B
Gemma 8B
Table 2: LLMs selected for each language. The best performing LLM for each language is in bold.
Prompt Selection: We started with  Silva et al. (2023)’s prompt, which was originally designed for agriculture examination contexts. However, our long-term goal is to support farmers with advisories. Hence, we adapted that prompt and generated six different prompt variations, which are discussed in Appendix D. Zero-shot experiments were conducted to select the best prompt.
No. Parameter Evaluation Question Scale
Table 3: Evaluation criteria for human assessment
LLM Fine-tuning: The best performing LLM under the zero-shot setting was fine-tuned with the best prompt, using the synthetic dataset we created. We also conducted an experiment to evaluate the efficacy of the translate-test approach for Hindi and Punjabi. Implementation details and hyperparameters are in Appendix E.
3.2 Evaluation Metrics
LLM response to each question was evaluated by human evaluators using three metrics (see Table 3): Agricultural Consensus, Relevancy, and Factuality. These metrics are inspired by similar evaluation methods used in prior research Singhal et al. (2023); Balaguer et al. (2024). Human evaluator details and evaluation guidelines are outlined in Appendix  F.
Setting LLMs/ Temp Unique Total
Prompts QA pairs Evals
Multilingual LLM Comparison
English 4 LLMs 3 20 720*
Hindi 5 LLMs 1 100 1500*
Punjabi 3 LLMs 1 100 900*
Prompt Selection (1 LLM)
English 6 Prompts 1 20 120
Base vs. Finetuned Evaluation
English 2 (Base+FT-EN) 1 300 600
Hindi 2 (Base+FT-HI) 1 240 480
Punjabi 2 (Base+FT-PA) 1 240 480
Translate Test Approach (via FT-EN model)
Hindi 1 (FT-EN) 1 240 240
Punjabi 1 (FT-EN) 1 240 240
Total Evaluations 5,280
Table 4: Number of QA samples used for human experiments. Temperature values 0.2, 0.5, 0.7 were tested only for English model selection.*Each QA pair was independently evaluated by 3 annotators.
A subset of the QA pairs (see Table 4) was evaluated by three annotators. Multiple raters/measurements Koo et al. (2016) was used to calculate inter-annotator agreement for relevancy and factuality. Fleiss Kappa  Fleiss et al. (1971) was used for agriculture consensus. Why these measures were selected and how to interpret the results are in Appendix G. Inter-rater agreement is reported in Table 5. Agreement for consensus is almost perfect, while agreement for relevancy and factuality show moderate to good agreement.
4 Results
Due to the scarcity of human evaluators, it was not possible to evaluate all the LLMs across the full test set. Therefore, a subset of QA pairs was selected from the test set. From each QA pair, the question was used in the prompt, and the corresponding answer generated by the LLM was evaluated against the ground-truth answer by a human. Table 4 shows the number of human evaluations carried out.
4.1 Model and Prompt Selection
Instruction:
Provide only the answer relevant to the question. Do not include any other information. Answer should be in the same language as the question.
Input:
Question: {input}
Response:
Figure 2: Prompt used for answer generation for a given question.
Among the six tested prompts, Prompt 1 (see Figure 2) achieved the highest scores for English with 100%, 95.0% and 55.0% in consensus, relevancy and factuality (respectively) and was therefore used in the subsequent experiments. Results for all the prompts are in Table 15 of Appendix H.
Results in Table 6 show that Llama-3-8B-Instruct, Llama-3.1-8B-Instruct and Llama-3.1-8B-Instruct emerged as the best LLMs for English, Hindi and Punjabi, respectively, at temperature T = 0.2. Experiments were also conducted at T =0.5 and 0.7 for English (results are in Appendix H).
Metric Calculation English Hindi Punjabi
Consensus
Fleiss’ $\kappa$ 1.0 1.0 0.95
Relevancy ICC 0.79 0.62 0.77
Factuality ICC 0.82 0.93 0.75
Table 5: Inter-rater agreement for the three languages.
Lang. LLM Con.(%) Rel.%) Fact.(%)
English
Airavata-7B-Instruct 100.0 58.3 30.0
DeepSeek-8B-Instruct 100.0 33.3 28.3
Llama-3-8B-Instruct 100.0 98.3 66.6
Mistral-7B-Instruct 98.3 93.3 30.0
Hindi
Gemma-7B-Instruct 90 26.3 14.6
Llama-3-8B-Instruct 84.0 25.3 13.3
Llama-3.1-8B-Instruct 96.0 74.6 35.6
Mistral-7B-Instruct 100.0 37.3 5.6
Navarasa-7B-Instruct 91.0 58.6 21.0
Punjabi
Llama-3-8B-Instruct 87.0 17.6 12.3
Llama-3.1-8B-Instruct 100.0 43.3 24.3
Navarasa-7B-Instruct 95.6 29.3 15.6
4.2 Base vs. Fine-Tuned LLM Results
For each language, the LLM that provided the best result in the zero-shot setup was used for fine-tuning experiments. Results are in Table 7. A notable improvement was observed for relevancy and factuality across all languages after fine-tuning, despite a drop in performance for consensus for Hindi and Punjabi. The results also indicate that the English fine-tuned LLM in translate-test approach outperformed the Hindi and Punjabi fine-tuned LLMs. Results are the best for English, followed by Hindi and Punjabi. This could be a result of differing data amounts used in LLM pre-training.
Lang Consensus(%) Relevancy(%) Factuality(%)
BASE FT TT BASE FT TT BASE FT TT
En 98.0 99.3 32.3 70.6 32.6 55.3
Hi 97.5 93.3 96.2 62.5 66.6 74.1 28.7 38.3 42.0
Pa 95.0 85.4 96.6 38.3 62.5 72.9 25.8 35.4 37.9
Table 7: Comparison of Base, Fine-Tuned (FT) and Translate-test with FT-EN LLM (TT) performance.
5 Challenges with LLM-as-a-Judge Evaluation
In the above experiments, 5,280 human evaluations were done. Although human evaluations are the gold standard for assessing LLM performance, they are time consuming and expensive Tan et al. (2025). As a solution, we employed the "LLM-as-a-Judge" approach for evaluation. We used three instruction-tuned LLMs, namely Llama-4-Scout-17B-16E-Instruct, Llama-4-Maverick-17B-128E-Instruct, and DeepSeek-R1-Distill-Llama-70B [^3] We excluded larger LLMs like Llama 4 Behemoth (288B) and DeepSeek-R1 (685B) due to resource constraints.. We generated prompts, each corresponding to the evaluation metrics ie. consensus, factuality and relevancy (See Appendix  I for prompts), which were adapted from G-eval  (Liu et al., 2023). The same QA pairs used for human evaluation were evaluated using LLM judges.
As it can be observed in an example given in Table 8 (More examples in Appendix L.), the LLM judge misclassifies LLM response as partially correct. Bertscore, ROUGE and BLEU values are inconsistent, making it difficult to reach a judgment. Correlation between LLM judges and humans is shown in Table 9. Humans and the LLM Judges had a near-perfect agreement for consensus, thus are not reported. However, even the best result produced by the LLM judges showed a week or moderate correlation for both relevance and factuality, questioning the validity of LLM-as-Judge for domain, language and country-specific QA tasks.
Question:
What is the cause for round scales on the under surface of leaves along with yellowing of leaves on the upper surface in coconut palm?
LLM Answer:
Scale insects reside on the under surface of leaves and cause yellowing on the upper surface of leaves. They secrete wax in circular patches.
Gold Answer:
Fungal infection (Colletotrichum gloeosporioides) and insect infestation (Necrobia rufipes) are the common causes of round scales on the under surface of leaves along with yellowing of leaves on the upper surface in coconut palm.
Evaluation:
Expert Evaluation (Factuality): Incorrect LLM-as-Judge Evaluation (Factuality): Partially Correct BERT F1 Score: 0.6710 , ROUGE-L: 0.4482 , BLEU: 0.2216
Table 8: Example result by automated evaluation
Lang. LLM Rel. (Spearman $\rho$) Fact. (Spearman $\rho$)
En DeepSeek 0.23 0.27
L4-Scout 0.18 0.46
L4-Maverick 0.24 0.37
Hi DeepSeek 0.49 0.44
L4-Scout 0.45 0.49
L4-Maverick 0.49 0.45
Pa DeepSeek 0.35 0.23
L4-Scout 0.40 0.39
L4-Maverick 0.39 0.38
Table 9: Correlation between LLMs and humans.
6 Conclusion
In this work, we introduced a synthetic fine-tuning dataset as well as a human-curated evaluation dataset in English, Hindi, and Punjabi for question-answering in the Indian agricultural context. Our experiments demonstrate that fine-tuned LLMs significantly outperform their base counterparts, underscoring the value of synthetic data. Furthermore, for Hindi and Punjabi, the translate-test approach yields better performance than directly fine-tuning LLMs in those languages. As the next step, we plan to further validate our results through field testing with farmers. We also plan to explore techniques such as RAG or LLM agents to get up-to-date information on current weather, crop prices, etc.
7 Limitations
There are several limitations to our work. First, the dataset created and the methodology used in this study include only textual data. However, in real-world scenarios, farmers may have image-based queries such as pictures of crops or pests/diseases. It is not addressed by our current approach. Our future work will focus on building a multimodal and multilingual dataset that can handle both textual and visual input. It will also incorporate a wider range of Indian languages. Second, human evaluation was done only on a subset of the dataset, due to the resource-intensive nature of the task. In future, a more comprehensive human evaluation across the entire dataset is necessary to ensure LLM’s efficacy and deployment. Another limitation is that is that our experiments were conducted using relatively smaller LLMs (7B to 8B parameter range) due to limited computational resources. In future, the effect of using larger LLMs on performance can be explored. Our current work does not account for dynamic contextual factors such as weather conditions or soil health, which are often important in generating accurate and dynamic advisories. Integrating such contextual data through various APIs in an agentic framework can enable adaptive and real-time responses and it is an important direction for our future work.
8 Ethical Considerations
The human annotators volunteered their time to contribute to the project, and no financial compensation was provided. Our approach follows the principles of participatory research Nekoto et al. (2020); Gamage et al. (2025), and data creators have been offered coauthorship to compensate for their contributions. We crawled the data by adhering to the robots.txt specifications for each website.
Appendix A Data Source
@@
Appendix B Evaluation Dataset Creation
The evaluation dataset used in this study for English, Hindi and Punjabi languages was curated by Kashish Mahendra, Vidhi and Sudhir Rajput. It was manually extracted and verified from the websites given in the Table 13
Source Weblink
Tamil Nadu Agricultural University http://www.agritech.tnau.ac.in/expert_system/
Union Bank of India https://www.unionbankofindia.co.in/pdf/faq-agri-products.pdf
ICAR-NRRI, Cuttack https://icar-nrri.in/faq/
ICAR-NRCP, Solapur https://nrcpomegranate.icar.gov.in/FAQ
DD Kisan YouTube Channel https://www.youtube.com/watch?v=MYkBswpxRcA
Table 11: Sources and their corresponding web links used for evaluation dataset creation.
Appendix C LLM Details
LLM Name Size Section(s) Used
Llama-3-8B-Instruct 8.03B LLM Selection (EN, HI, PA) Fine-tuning (EN) Prompt Selection
Mistral-7B-Instruct-v0.3 7.25B LLM Selection (EN, HI)
Llama-3.1-8B-Instruct 8.03B LLM Selection (HI, PA) Fine-tuning (HI, PA)
Indic-gemma-7b-finetuned-sft-Navarasa 7B LLM Selection (HI, PA)
DeepSeek-R1-Distill-Llama-8B 8.03B LLM Selection (EN)
Airavata-7B-Instruct 6.87B LLM Selection (EN)
gemma-7b-it 7B LLM Selection (HI)
DeepSeek-R1-Distill-Llama-70B 70.6B Synthetic QA Generation LLM-as-a-Judge
Llama-4-Scout-17B-16E-Instruct 109B LLM-as-a-Judge
Llama-4-Maverick-17B-128E-Instruct 402B LLM-as-a-Judge
Table 12: List of LLMs used in the study, their parameter sizes, and the sections they were used in. HI: Hindi, PA: Punjabi, EN: English.
Appendix D Prompt Selection
This section provides the six tested prompts used for evaluating Mistral-7B in the Prompt Selection experiment and the design intuition behind them. Each prompt was designed to optimize the quality of the response for agricultural question-answer generation.
Prompts Tested